Determine whether the system of linear equations has one and only one solution, infinitely many solutions, or no solution.
2x − y= −5
4x − 2y = 6
Is there one and only one solution?
infinitely many solutions ? Or no solution?
Find the solution, if one exists.(If there are infinitely many solutions, express x and y in terms of the parameter t. If there is no solution, enter NO SOLUTION.)

Answer by ewatrrr(24785)


Hi
2x − y= −5 |multiplying thru by -2 to eliminate the y variable by adding the Eqs
4x − 2y = 6
-4x + 2y= 10
4x − 2y = 6
0 ≠ 4 There is NO Solution